OpenStudy (anonymous):

The solution of x + 5/2 = 15 is 25. True or False?

jimthompson5910 (jim_thompson5910):

If the equation is \[\Large \frac{x+5}{2} = 15\] then \[\Large \frac{x+5}{2} = 15\] \[\Large \frac{25+5}{2} = 15\] \[\Large \frac{30}{2} = 15\] \[\Large 15 = 15\] which is true, so x = 25 is indeed the solution

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome